OpenStudy (anonymous):

The equation of a surface in spherical coordinates is rho^2(1+2cos^2(phi))=1 Identify this surface; what is its Cartesian equation

OpenStudy (anonymous):

I know that we have to convert from spherical to Retangular or Cartesian and the transformati9on equations are : rho= sqrt(x^2+y^2+z^2) tan theta = y/x cos phi = z/p But i don't know how to apply it to the equation given because pieces seem to be missing